# Triple DES AES Advanced Encryption Standard

### v2

@@ 4.2. Discussion
Regarding GPU memory protection, many designs assume that the adversary cannot access secrets in the GPU memory or explicitly state that physical attacks targeting memory are out of their scope. For instance, Graviton (Volos et al., 2018) and Honeycomb (Mai et al., 2023) trust the device memory of the GPU, as modern GPUs typically integrate device memory using 2.5D/3D silicon interposers within the same package. These designs mainly focus on preventing adversaries from accessing sensitive data stored in GPU memory with finer-grained access control. To fill this gap, LITE (Yudha et al., 2022) designs a software-based unified encryption domain to protect sensitive data stored in GPU memory, including L1/L2 cache and device memory. Additionally, NVIDIA H100 (NVIDIA, 2022a) and TNPU (Lee et al., 2022) support hardware-based memory encryption, while CAGE (Wang et al., 2024b) requires hardware-assisted memory encryption support built into future CCA devices.
@@
Hardware changes. Earlier works like Graviton (Volos et al., 2018) and HIX (Jang et al., 2019) require non-trivial modifications to peripheral components, such as the GPU’s command processor, the introduction of new SGX instructions, specific Memory Management Unit (MMU) page table walkers, and other hardware changes. The extent to which a GPU TEE design modifies hardware components determines its practical compatibility, i.e., its suitability for various GPUs. Many designs necessitate hardware changes that have long lead times before they can be deployed in production environments. Consequently, works like LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), and Honeycomb (Mai et al., 2023), seek purely software-based approaches for trusted GPU execution. CAGE (Wang et al., 2024b) primarily leverages generic hardware features in Arm CCA and GPUs without requiring additional hardware changes or customization of existing CPUs and GPUs.
@@
TCB size. One of the key challenges in GPU TEE design is reducing the TCB size as much as possible. A large software stack increases the TCB, which can introduce more vulnerabilities and hence threaten security. An effective and efficient way to reduce the TCB is to place most of the GPU software stack in the untrusted OS, while keeping only the key functions for controlling the GPU in a trusted area of the CPU, such as an SGX enclave. For example, CAGE (Wang et al., 2024b) achieves a ‘thin’ TCB for realms by using its shadow task mechanism instead of loading the heavyweight GPU software into the TEE. It introduces 2–-26K lines of code for a CPU-side isolation and 1,301 lines of code for monitor, resulting in the lowest TCB size among current Arm-based GPU TEEs.
@@
Performance overhead. The ultimate goal of designing a GPU TEE is to ensure strong isolation of GPU execution while minimizing the performance overhead associated with providing security guarantees. Comparing the performance overhead of different schemes is challenging due to the costly re-implementation efforts required, as each scheme targets different platforms and is implemented under varying conditions. For example, in HIX, the software components are implemented on top of an emulated system (like KVM and QEMU), which supports hardware modifications. Several factors, such as the number of cryptographic operations, the frequency of switching between secure and normal worlds, and the integrity checks, play significant roles in optimizing performance compared to insecure baselines on GPUs. For example, CAGE (Wang et al., 2024b) outperforms StrongBox (Deng et al., 2022) by 1.96%–13.16% on six Rodinia benchmarks. This performance gain is attributed to CAGE’s ability to securely transfer plaintext data to buffers and store the results without introducing additional cryptographic operations. It is reasonable to speculate that the NVIDIA H100 GPU introduces a smaller performance burden due to its hardware-based encryption support. It would be insightful to conduct a comprehensive performance comparison across different platforms once the entire workflow for confidential computing on the NVIDIA H100 GPU is fully available in the future.
@@ 5. Security of GPUs
Due to their widespread adoption as accelerators, GPUs have come under scrutiny from both offensive and defensive perspectives in recent years. Researchers have explored GPU vulnerabilities at hardware architecture and software levels, showcasing practical attacks and suggesting mitigations. We believe that summarizing and analyzing attacks on GPUs will facilitate the exploration of potential attacks on GPU TEEs. While Naghibijouybari et al. (Naghibijouybari et al., 2022) offer a comprehensive survey categorizing GPU security vulnerabilities and potential countermeasures, their analysis does not cover significant advancements that have appeared since the latter half of 2022. Moreover, their emphasis primarily is on microarchitectural attacks across diverse accelerators without considering attacks that leverage physical access or measurements. Nonetheless, the consideration of physical channels is essential for potential attacks on GPU TEE. In this section, we review existing GPU attacks, including microarchitectural side-channel/covert-channel attacks, physical side-channel attacks, and fault injection attacks. LABEL:tab:GPU_attacks provides a summary of attacks, target GPUs, threat model, attack goals and methods.
{ThreePartTable}{TableNotes}
CNN/DNN denotes Convolutional and Deep Neural Network. In Attacks, SCA, CovA, and FIA denote Side-Channel Analysis, Covert-Channel, and Fault Injection Attacks, respectively. AA/LA represents an attack that exploits an architectural or logical vulnerability on GPU. Physical denotes whether this attack leverages physical properties (e.g., power consumption, EM, acoustics, etc.). In Model, sGPU, mGPU, and iGPU denote the attack is deployed on a single GPU, multiple GPUs or integrated GPU, respectively. Note that multiple GPUs represent multiple discrete GPU cards or multiple GPU instances (e.g., NVIDIA’s MIG feature). In Main results, we report the central outcomes, e.g., bandwidth and classification accuracy for CovA and SCA, respectively.
@@
Work&Year Accelerator Attack Threat Model Phy- sical Attack Scenarios & Goals Method Main results
@@
Lee et al.’14 (Lee et al., 2014) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Infer which web pages a victim user has visited Exploit that GPUs do not initialize newly allocated memory pages $95.4\%$ accuracy
@@
Maurice et al.’14 (Maurice et al., 2014) NVIDIA GPU AA/LA sGPU Require root privileges based on accessing GPU memory via GPU runtime or PCI configuration Recover data of victim’s previously execution on GPU in virtualized environment Exploit the vulnerability of memory clearing on GPU Access data stored in GPU global memory after soft reboot
Luo et al.’15 (Luo et al., 2015) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover all of the last round Advanced Encryption Standard (AES) key byte Apply correlation power analysis to analyze the power consumption of the GPU Recover the key with 160,000 power traces
@@
Pietro et al.’16 (Pietro et al., 2016) NVIDIA GPU AA/LA sGPU User-level permissions Read information stored in memory and infer the AES keys Leverage the lack of adequate data allocation and data cleaning $70\%$, $28\%$ leakages
@@
Zhou et al.’16 (Zhou et al., 2016) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Recover users’ profile images and email contents Reconstruct images from identified image-like tiles in GPU memory $79.3\%$ successful recovery
Jiang et al.’16 (Jiang et al., 2016) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Leverage the relation between encryption kernel’s runtime and the key 30 min with 1M timing samples
Jiang et al.’17 (Jiang et al., 2017) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Exploit timing variability due to shared memory bank conflicts. Use 10M timing samples
Naghibijouybari et al.’17 (Naghibijouybari et al., 2017) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permission Infer transmitted data Reverse engineer warp schedulers; learn contention on caches and memory $4$ MB/s (error-free)
Frigo et al.’18 (Frigo et al., 2018) Intel HD, Qualcomm Adreno SCA, FIA iGPU Access iGPU through a malicious application or when victim visits malicious JavaScript website Escape the Firefox sandbox on Android platforms and break ASLR Detect contiguous memory allocation via SCA and then trigger Rowhammer bit flips Hijack the browser and break ASLR in 116 s and 27 s
@@
Gao et al.’18 (Gao et al., 2018) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover AES private key Sample MMIO traces via trigger; exploit parallel MMIO leakages Recover the key with 11k MMIO traces
Naghibijouybari et al.’18 (Naghibijouybari et al., 2018) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Infer victim’s accessed website data and the number of neurons Learn contention via memory allocation, performance counter, and timing $90\%$, $94.8\%$, $88.6\%$ accuracy
Luo et al.’19 (Luo et al., 2019) NVIDIA GPU SCA sGPU User-level permissions; available to timing information Recover RSA private key Capture the parallel characteristics of an RSA on GPU via timing model Recover 512/2048-bit key in 15 min/8 h
Sabbagh et al.’20 (Sabbagh et al., 2020) AMD GPU FIA sGPU DVFS interfaces access Recover AES key Exploit DVFS to introduce random faults during kernel execution Recover key in 15 min
Wei et al.’20 (Wei et al., 2020) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Recover the layer composition and hyper-parameters of DNN Leverage context-switching penalties when switching different kernels $76.9\%$-$100\%$ accuracy
Hu et al.’20 (Hu et al., 2020) NVIDIA GPU SCA sGPU Physically access the victim GPU with user-level permissions $\checkmark$ Obtain the complete model architecture information Learn the inter-layer temporal association information from kernel execution $75.9\%$ accuracy
HE et al.’20 (HE et al., 2020) Intel HD 530 AA/LA, CovA iGPU Unprivileged adversary and the victim co-locate on the same GPU, but are isolated (1) Recover AES key; (2) monitor the browser activity; (3) conduct a covert channel to exchange data Exploit context management with uncleared registers and shared memory during context switches Recover AES-128/192 in 0.15 s/2 min; $90.7\%$ accuracy; 1.3–8 Gbps
Santos et al.’21 (Santos et al., 2021) NVIDIA GPU FIA sGPU Inject faults into pipeline registers, ALU, and memory Improve the probability of incorrect output by injecting faults Combine register transfer level FIA with software FIA Underestimate the error rate by up to $48\%$
Ahn et al.’21 (Ahn et al., 2021b) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permissions Infer transmitted data Exploit interconnect channel contention $24$ MB/s
Cronin et al.’21 (Cronin et al., 2021) Apple GPU SCA sGPU Exploit JavaScript website and trick user into installing a malicious app Track the user’s web activity Exploit the shared cache between the CPU-cores and peripherals $90\%$, $95\%$ accuracy
Zhu et al.’21 (Zhu et al., 2021) NVIDIA GPU Snooping sGPU Install a PCIe bus snooping device to monitor and log the PCIe traffic Recover the DNN model (semantically identical architecture) Reverse engineer GPU command headers; collect PCIe packets via snooping Similar accuracy as original ones
@@
Ahn et al.’21 (Ahn et al., 2021a) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Recover early/later key bytes via timing correlation and cache-collision Use 100k timing samples
@@
Dutta et al.’21 (Dutta et al., 2021) Intel HD Graphics CovA iGPU Spy and trojan run on CPU and GPU with user-level permissions Infer transmitted data and track user’s cache activity CovA1: prime+probe attack on LLC; CovA2: observe delays on the ring bus CovA1: 120 KB/s, CovA2: 400 KB/s
@@
Nayak et al.’21 (Nayak et al., 2021) NVIDIA GPU CovA sGPU User-level permissions Infer the accessed data from Virginian database Leverage the shared last-level TLB (reverse engineering) and prime+probe 81 KB/s (MPS-enabled)
@@
Yang et al.’22 (Yang et al., 2022) Qualcomm Adreno SCA sGPU Install and launch an eavesdropping application on the victim device Eavesdrop the user’s input credentials (login usernames and passwords) through onscreen keyboard Read raw values from the GPU device file and identify GPU events from key presses via ML models Correctly infer more than $80\%$ of user’s credential inputs
@@
Zhan et al.’22 (Zhan et al., 2022) NVIDIA/AMD GPU SCA sGPU Unprivileged attacker eavesdrops on a victim $\checkmark$ Website fingerprinting and keystroke timing inference Exploit EM caused by DVFS $95.6\%$ accuracy and $90\%$ keystrokes
@@
Liang et al.’22 (Liang et al., 2022) NVIDIA GPU SCA sGPU Attacker and victim GPUs may be separated by meters $\checkmark$ Infer the number of layers/kernels, sizes of layers and strides Exploit far-field EM signals emanated from a GPU Recover VGG’s and ResNet’s parameters
@@
Maia et al.’22 (Maia et al., 2022) NVIDIA GPU SCA sGPU Unprivileged attacker with physical proximity $\checkmark$ Infer neural network architectures and layer parameters Exploit the magnetic flux emanating from a GPU’s power cable Up to $96.8\%$ accuracy
@@
Horvath et al.’23 (Horvath et al., 2023) NVIDIA GPU SCA sGPU User-level permission with physical access and know the CNN architecture $\checkmark$ Recover the weights and biases of CNN Correlation electromagnetic analysis N/A
@@
Dutta et al.’23 (Dutta et al., 2023) NVIDIA GPU SCA, CovA mGPU Spy and victim connect via NVLink with user-level permissions CovA: infer message; SCA: infer the number of neurons of NN Mount Prime+Probe attack on the L2 cache of victim GPU CovA: 4 MB/s; SCA: $99.91\%$ accuracy
@@
Zhang et al.’23 (Zhang et al., 2023) NVIDIA GPU CovA mGPU Unprivileged spy and trojan on an MIG-enabled GPU Infer transmitted data and classify the running ML frameworks Mount Prime+Probe attack on the shared L3-TLB $31.47$ KB/s
@@
Taneja et al.’23 (Taneja et al., 2023) Various GPUs SCA iGPU, sGPU Read frequency, power, and temperature from internal sensors (may require privileged access) $\checkmark$ (1) Observe CPU/GPU behaviour; (2) recover images and user’s browsing history; (3) fingerprint websites (1) Use data from internal sensors; (2) observe throttling via timing; (3) leverage frequency and power bursts on iGPUs Pixel stealing with accuracy of 60%–90%
@@
Wang et al.’24 (Wang et al., 2024a) Various GPUs SCA iGPU, sGPU Observe coarse-grained pattern redundancy using timer Perform cross-origin pixel stealing attacks in the browser Infer data-dependent DRAM traffic and cache usage via compression channel Pixel stealing with accuracy 68.6%–99.6%
@@
Giner et al.’24 (Giner et al., 2024) NVIDIA/AMD GPU SCA, CovA sGPU SCA: Victim passively visits attacker’s webpage; CovA: Spy accesses the website visited by victim SCA: Detect typed characters and recover the victim’s AES key; CovA: Infer transmitted data Exploit cache contention on the L2 cache via Prime+Probe attack, parallelize eviction set construction SCA: $98\%$ accuracy, recover key in 6  min; CovA: $10.9$ KB/s
[…]
5.1. Architectural/Logical Attacks

### v3

@@ 4.2. Discussion
Comparison to CPU TEEs. While both CPU and GPU TEEs aim to provide secure and isolated execution environments, their designs differ significantly due to fundamental differences in microarchitecture and usage patterns. CPU TEEs, such as Intel SGX and AMD SEV-SNP, operate in control-centric environments with centralized execution and predictable memory behavior, making them well-suited for fine-grained access control, secure paging, and privilege enforcement. In contrast, GPUs are optimized for massive data-parallel workloads, with thousands of lightweight threads, wide memory bandwidth, and deep memory hierarchies. These characteristics make it difficult to directly apply traditional TEE techniques like enclave management or synchronized memory encryption. Moreover, GPUs often lack built-in privilege separation and rely on the CPU for task scheduling and memory control, expanding the attack surface. As a result, GPU TEE designs typically require either hardware modifications to decouple from the host (e.g., Graviton, HIX) or leverage CPU-side TEEs for protection and communication (e.g., Honeycomb, SAGE). This architectural gap explains the greater diversity and complexity in GPU TEE designs compared to their CPU counterparts.
@@
Regarding GPU memory protection, many designs assume that the adversary cannot access secrets in the GPU memory or explicitly state that physical attacks targeting memory are out of their scope. For instance, Graviton (Volos et al., 2018) and Honeycomb (Mai et al., 2023) trust the device memory of the GPU, as modern GPUs typically integrate device memory using 2.5D/3D silicon interposers within the same package. These designs mainly focus on preventing adversaries from accessing sensitive data stored in GPU memory with finer-grained access control. To fill this gap, LITE (Yudha et al., 2022) designs a software-based unified encryption domain to protect sensitive data stored in GPU memory, including L1/L2 cache and device memory. Additionally, NVIDIA H100 (NVIDIA, 2022a) and TNPU (Lee et al., 2022) support hardware-based memory encryption, while CAGE (Wang et al., 2024a) requires hardware-assisted memory encryption support built into future CCA devices.
@@
Hardware changes. Earlier works like Graviton (Volos et al., 2018) and HIX (Jang et al., 2019) require non-trivial modifications to peripheral components, such as the GPU’s command processor, the introduction of new SGX instructions, specific Memory Management Unit (MMU) page table walkers, and other hardware changes. The extent to which a GPU TEE design modifies hardware components determines its practical compatibility, i.e., its suitability for various GPUs. Many designs necessitate hardware changes that have long lead times before they can be deployed in production environments. Consequently, works like LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), and Honeycomb (Mai et al., 2023), seek purely software-based approaches for trusted GPU execution. CAGE (Wang et al., 2024a) primarily leverages generic hardware features in Arm CCA and GPUs without requiring additional hardware changes or customization of existing CPUs and GPUs.
@@
TCB size. One of the key challenges in GPU TEE design is reducing the TCB size as much as possible. A large software stack increases the TCB, which can introduce more vulnerabilities and hence threaten security. An effective and efficient way to reduce the TCB is to place most of the GPU software stack in the untrusted OS, while keeping only the key functions for controlling the GPU in a trusted area of the CPU, such as an SGX enclave. For example, CAGE (Wang et al., 2024a) achieves a ‘thin’ TCB for realms by using its shadow task mechanism instead of loading the heavyweight GPU software into the TEE. It introduces 2–-26K lines of code for a CPU-side isolation and 1,301 lines of code for monitor, resulting in the lowest TCB size among current Arm-based GPU TEEs.
@@
Performance overhead. The ultimate goal of designing a GPU TEE is to ensure strong isolation of GPU execution while minimizing the performance overhead associated with providing security guarantees. Comparing the performance overhead of different schemes is challenging due to the costly re-implementation efforts required, as each scheme targets different platforms and is implemented under varying conditions. For example, in HIX, the software components are implemented on top of an emulated system (like KVM and QEMU), which supports hardware modifications. Several factors, such as the number of cryptographic operations, the frequency of switching between secure and normal worlds, and the integrity checks, play significant roles in optimizing performance compared to insecure baselines on GPUs. For example, CAGE (Wang et al., 2024a) outperforms StrongBox (Deng et al., 2022) by 1.96%–13.16% on six Rodinia benchmarks. This performance gain is attributed to CAGE’s ability to securely transfer plaintext data to buffers and store the results without introducing additional cryptographic operations. It is reasonable to speculate that the NVIDIA H100 GPU introduces a smaller performance burden due to its hardware-based encryption support. It would be insightful to conduct a comprehensive performance comparison across different platforms once the entire workflow for confidential computing on the NVIDIA H100 GPU is fully available in the future.
Trade-off between cryptographic methods and GPU TEEs. In general, while cryptographic approaches like HE and MPC offer strong confidentiality guarantees, they remain impractical for many GPU workloads due to significant computation/communication overhead, memory usage, and developer complexity. Recent GPU-based MPC systems such as CryptGPU (Tan et al., 2021), Piranha (Watson et al., 2022), and Orca (Jawalkar et al., 2024) show that performance is highly sensitive to the number of computing participants and choice of operators, making them difficult to develop. Piranha further highlights that communication overhead becomes a major bottleneck in wide-area network settings—an important consideration since real-world deployments often involve geographically distributed GPU providers. For HE, even with dedicated HE accelerator support, as in TensorFHE (Fan et al., 2023), significant modifications are needed to underlying schemes (e.g., CKKS) and operations (e.g., NTT). Applying HE to ML workloads also requires careful tuning of HE parameters and may necessitate changes to model architecture, complicating deployment in applications. Therefore, GPU TEEs offer a more practical path forward, enabling protection of existing workloads with less developer effort and minimal modification to ML models. While GPU TEEs are still in their early stages and only applicable on certain machines, we believe they hold strong potential for future GPU-based applications, especially in ML.
@@ 5. Security of GPUs
Due to their widespread adoption as accelerators, GPUs have come under scrutiny from both offensive and defensive perspectives in recent years. Researchers have explored GPU vulnerabilities at hardware architecture and software levels, showcasing practical attacks and suggesting mitigations. We believe that summarizing and analyzing attacks on GPUs will facilitate the exploration of potential attacks on GPU TEEs. While Naghibijouybari et al. (Naghibijouybari et al., 2022) offer a comprehensive survey categorizing GPU security vulnerabilities and potential countermeasures, their analysis does not cover significant advancements that have appeared since the latter half of 2022. Moreover, their emphasis primarily is on microarchitectural attacks across diverse accelerators without considering attacks that leverage physical access or measurements. Nonetheless, the consideration of physical channels is essential for potential attacks on GPU TEE.
In this section, we review existing GPU attacks, including microarchitectural side-channel/covert-channel attacks, physical side-channel attacks, and fault injection attacks. LABEL:tab:GPU_attacks provides a summary of attacks, target GPUs, threat model, attack goals and methods. Based on LABEL:tab:GPU_attacks, we highlight several interesting take-away points: ❶ Most attacks target single discrete GPU systems, with memory and timing behavior serving as the primary leakage vectors. However, recent research has begun to explore attacks on multi-GPU and multi-instance GPU environments, revealing new vulnerabilities in GPU-to-GPU communication and resource partitioning. ❷ A majority of attacks can be carried out with only user-level permissions, indicating that isolation mechanisms in modern GPU systems remain inadequate, particularly in multi-tenant or cloud-based deployments. ❸ Attacks span across multiple GPU vendors (e.g., NVIDIA, AMD, Qualcomm) and consistently aim to recover sensitive information, such as AES keys, Neural Network (NN) model parameters, and user input keystrokes. Notably, attacks targeting more complex GPU workloads, such as Large Language Models, remain largely unexplored. ❹ While impactful, fault injection and covert channels appear less frequently in the literature. These often require stronger attacker capabilities or intricate coordination between software and hardware layers, which may explain their limited exploration. ❺ Emerging attacks have begun to explore cross-GPU timing channels and remote leakage via browsers, suggesting that cloud-hosted GPU services—especially those accessible by multiple users—may become an increasingly important and vulnerable attack surface.
@@
Work&Year Accelerator Threat Model Phy- sical Attack Scenarios & Goals Method Main results
@@
Lee et al.’14 (Lee et al., 2014) NVIDIA/AMD GPU sGPU User-level permissions Infer which web pages a victim user has visited Exploit that GPUs do not initialize newly allocated memory pages $95.4\%$ accuracy
@@
Maurice et al.’14 (Maurice et al., 2014) NVIDIA GPU sGPU Require root privileges based on accessing GPU memory via GPU runtime or PCI configuration Recover data of victim’s previously execution on GPU in virtualized environment Exploit the vulnerability of memory clearing on GPU Access data stored in GPU global memory after soft reboot
@@
Pietro et al.’16 (Pietro et al., 2016) NVIDIA GPU sGPU User-level permissions Read information stored in memory and infer the Advanced Encryption Standard (AES) keys Leverage the lack of adequate data allocation and data cleaning $70\%$, $28\%$ leakages
@@
Zhou et al.’16 (Zhou et al., 2016) NVIDIA/AMD GPU sGPU User-level permissions Recover users’ profile images and email contents Reconstruct images from identified image-like tiles in GPU memory $79.3\%$ successful recovery
Microarchitectural Side-Channel/Covert-Channel Attacks
@@
Luo et al.’15 (Luo et al., 2015) NVIDIA GPU sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover all of the last round AES key byte Apply correlation power analysis to analyze the power consumption of the GPU Recover the key with 160,000 power traces
@@
Jiang et al.’16 (Jiang et al., 2016) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Leverage the relation between encryption kernel’s runtime and the key 30 min with 1M timing samples
Jiang et al.’17 (Jiang et al., 2017) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Exploit timing variability due to shared memory bank conflicts. Use 10M timing samples
Naghibijouybari et al.’17 (Naghibijouybari et al., 2017) NVIDIA GPU sGPU Trojan and spy co-exist on a GPU, both with user-level permission Infer transmitted data Reverse engineer warp schedulers; learn contention on caches and memory 44 MB/s (error-free)
Gao et al.’18 (Gao et al., 2018) NVIDIA GPU sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover AES private key Sample MMIO traces via trigger; exploit parallel MMIO leakages Recover the key with 11k MMIO traces
Naghibijouybari et al.’18 (Naghibijouybari et al., 2018) NVIDIA GPU sGPU Spy co-locates with the victim program with user-level permission Infer victim’s accessed website data and the number of neurons Learn contention via memory allocation, performance counter, and timing $90\%$, $94.8\%$, $88.6\%$ accuracy
Luo et al.’19 (Luo et al., 2019) NVIDIA GPU sGPU User-level permissions; available to timing information Recover RSA private key Capture the parallel characteristics of an RSA on GPU via timing model Recover 512/2048-bit key in 15 min/8 h
Wei et al.’20 (Wei et al., 2020) NVIDIA GPU sGPU Spy co-locates with the victim program with user-level permission Recover the layer composition and hyper-parameters of DNN Leverage context-switching penalties when switching different kernels $76.9\%$-$100\%$ accuracy
Hu et al.’20 (Hu et al., 2020) NVIDIA GPU sGPU Physically access the victim GPU with user-level permissions $\checkmark$ Obtain the complete model architecture information Learn the inter-layer temporal association information from kernel execution $75.9\%$ accuracy
HE et al.’20 (HE et al., 2020) Intel HD 530 iGPU Unprivileged adversary and the victim co-locate on the same GPU, but are isolated (1) Recover AES key; (2) monitor the browser activity; (3) conduct a covert channel to exchange data Exploit context management with uncleared registers and shared memory during context switches Recover AES-128/192 in 0.15 s/2 min; $90.7\%$ accuracy; 1.3–8 Gbps
Ahn et al.’21 (Ahn et al., 2021b) NVIDIA GPU sGPU Trojan and spy co-exist on a GPU, both with user-level permissions Infer transmitted data Exploit interconnect channel contention $24$ MB/s
Cronin et al.’21 (Cronin et al., 2021) Apple GPU sGPU Exploit JavaScript website and trick user into installing a malicious app Track the user’s web activity Exploit the shared cache between the CPU-cores and peripherals $90\%$, $95\%$ accuracy
Zhu et al.’21 (Zhu et al., 2021) NVIDIA GPU sGPU Install a PCIe bus snooping device to monitor and log the PCIe traffic Recover the DNN model (semantically identical architecture) Reverse engineer GPU command headers; collect PCIe packets via snooping Similar accuracy as original ones
Ahn et al.’21 (Ahn et al., 2021a) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Recover early/later key bytes via timing correlation and cache-collision Use 100k timing samples
@@
Dutta et al.’21 (Dutta et al., 2021) Intel HD Graphics iGPU Spy and trojan run on CPU and GPU with user-level permissions Infer transmitted data and track user’s cache activity CovA1: prime+probe attack on LLC; CovA2: observe delays on the ring bus CovA1: 120 KB/s, CovA2: 400 KB/s
@@
Nayak et al.’21 (Nayak et al., 2021) NVIDIA GPU sGPU User-level permissions Infer the accessed data from Virginian database Leverage the shared last-level TLB (reverse engineering) and prime+probe 81 KB/s (MPS-enabled)
@@
Yang et al.’22 (Yang et al., 2022) Qualcomm Adreno sGPU Install and launch an eavesdropping application on the victim device Eavesdrop the user’s input credentials (login usernames and passwords) through onscreen keyboard Read raw values from the GPU device file and identify GPU events from key presses via ML models Correctly infer more than $80\%$ of user’s credential inputs
@@
Zhan et al.’22 (Zhan et al., 2022) NVIDIA/AMD GPU sGPU Unprivileged attacker eavesdrops on a victim $\checkmark$ Website fingerprinting and keystroke timing inference Exploit EM caused by Dynamic Voltage and Frequency Scaling (DVFS) $95.6\%$ accuracy and $90\%$ keystrokes
@@
Liang et al.’22 (Liang et al., 2022) NVIDIA GPU sGPU Attacker and victim GPUs may be separated by meters $\checkmark$ Infer the number of layers/kernels, sizes of layers and strides Exploit far-field EM signals emanated from a GPU Recover VGG’s and ResNet’s parameters
@@
Maia et al.’22 (Maia et al., 2022) NVIDIA GPU sGPU Unprivileged attacker with physical proximity $\checkmark$ Infer NN architectures and layer parameters Exploit the magnetic flux emanating from a GPU’s power cable Up to $96.8\%$ accuracy
@@
Horvath et al.’23 (Horvath et al., 2023) NVIDIA GPU sGPU User-level permission with physical access and know the CNN architecture $\checkmark$ Recover the weights and biases of CNN Correlation electromagnetic analysis N/A
@@
Dutta et al.’23 (Dutta et al., 2023) NVIDIA GPU mGPU Spy and victim connect via NVLink with user-level permissions CovA: infer message; SCA: infer the number of neurons of NN Mount Prime+Probe attack on the L2 cache of victim GPU CovA: 4 MB/s; SCA: $99.91\%$ accuracy
@@
Zhang et al.’23 (Zhang et al., 2023) NVIDIA GPU mGPU Unprivileged spy and trojan on an MIG-enabled GPU Infer transmitted data and classify the running ML frameworks Mount Prime+Probe attack on the shared L3-TLB $31.47$ KB/s
@@
Taneja et al.’23 (Taneja et al., 2023) Various GPUs iGPU, sGPU Read frequency, power, and temperature from internal sensors (may require privileged access) $\checkmark$ (1) Observe CPU/GPU behaviour; (2) recover images and user’s browsing history; (3) fingerprint websites (1) Use data from internal sensors; (2) observe throttling via timing; (3) leverage frequency and power bursts on iGPUs Pixel stealing with accuracy of 60%–90%
@@
Wang et al.’24 (Wang et al., 2024b) Various GPUs iGPU, sGPU Observe coarse-grained pattern redundancy using timer Perform cross-origin pixel stealing attacks in the browser Infer data-dependent DRAM traffic and cache usage via compression channel Pixel stealing with accuracy 68.6%–99.6%
@@
Giner et al.’24 (Giner et al., 2024) NVIDIA/AMD GPU sGPU SCA: Victim passively visits attacker’s webpage; CovA: Spy accesses the website visited by victim SCA: Detect typed characters and recover the victim’s AES key; CovA: Infer transmitted data Exploit cache contention on the L2 cache via Prime+Probe attack, parallelize eviction set construction SCA: $98\%$ accuracy, recover key in 6  min; CovA: $10.9$ KB/s
Fault Injection Attacks
Frigo et al.’18 (Frigo et al., 2018) Intel HD, Qualcomm Adreno iGPU Access iGPU through a malicious application or when victim visits malicious JavaScript website Escape the Firefox sandbox on Android platforms and break ASLR Detect contiguous memory allocation via SCA and then trigger Rowhammer bit flips Hijack the browser and break ASLR in 116 s and 27 s
Sabbagh et al.’20 (Sabbagh et al., 2020) AMD GPU sGPU DVFS interfaces access Recover AES key Exploit DVFS to introduce random faults during kernel execution Recover key in 15 min
dos Santos et al.’21 (dos Santos et al., 2021) NVIDIA GPU sGPU Inject faults into pipeline registers, ALU, and memory Improve the probability of incorrect output by injecting faults Combine register transfer level FIA with software FIA Underestimate the error rate by up to $48\%$
CNN/DNN denotes Convolutional and Deep Neural Network. In Attacks, SCA, CovA, and FIA denote Side-Channel Analysis, Covert-Channel, and Fault Injection Attacks, respectively. AA represents an attack that exploits an architectural or logical vulnerability on GPU. Physical denotes whether this attack leverages physical properties (e.g., power consumption, EM, acoustics, etc.). In Model, sGPU, mGPU, and iGPU denote the attack is deployed on a single GPU, multiple GPUs or integrated GPU, respectively. Note that multiple GPUs represent multiple discrete GPU cards or multiple GPU instances (e.g., NVIDIA’s MIG feature). In Main results, we report the central outcomes, e.g., bandwidth and classification accuracy for CovA and SCA, respectively.
5.1. Architectural Attacks